# Complexity of Weighted First-Order Model Counting in the Two-Variable Fragment with Counting Quantifiers: A Bound to Beat

### v1

@@ 3.1 A Worked Example on Removing Counting Quantifiers
with $\overline{w}(D)=-1$. Apart from applying Lemma 2, we also introduced another fresh predicate $\overline{P}/1$ which wraps the negation of $P$ for brevity further down the line.
Next, we need to apply Lemma 5 three times, to Sentences 9 through 11. The repetitions can, luckily, be avoided. By the distributive property of conjunctions and disjunctions, we can factor $\overline{P}/1,C/1$ and $D/1$ out from the sentences, thus obtaining
@@
$\wedge\left(\forall x\;F(x)\Leftrightarrow\left(C(x)\vee D(x)\vee\overline{P}(x)\right)\right)$ (14)
@@
$\wedge\mathcal{T}_{l}(F,R)\wedge\left(\forall x\;F(x)\Leftrightarrow\left(C(x)\vee D(x)\vee\overline{P}(x)\right)\right)$
@@
where $B^{L}/2$ is a fresh predicate. We need Lemma 4 to get rid of this counting construct. Using said lemma will turn $\mathcal{T}_{l}(K,L)$ into
@@
$\wedge\mathcal{T}^{\prime\prime}_{l}(F,R)\wedge\left(\forall x\;F(x)\Leftrightarrow\left(C(x)\vee D(x)\vee\overline{P}(x)\right)\right)$
@@ 3.2 Deriving the Upper Bound
As one can observe, the formula $\varphi^{*}$ in the example above has grown considerably, compared to its original form. The question is whether the formula growth can influence the asymptotic bound from Theorem 1. At first glance, the answer may seem negative. That is due to the fact that the transformation only extends the vocabulary (which is assumed to be fixed), adds cardinality constraints (which are concentrated in the function $\alpha$) and increases length of the input formula (which is also a constant with respect to $n$). However, there is one caveat to be aware of. When extending the vocabulary, we may introduce new valid cells. The vocabulary is fixed once WFOMC computation starts, but if we have formula $\Gamma\wedge\Phi$ such that $\Gamma\in\textbf{FO}^{2}+\textbf{CC}{}$ and $\Phi\in\textbf{C}^{2}{}$, then this formula already has $p$ valid cells. Once we construct $\Gamma\wedge\Phi^{*}$, where $\Phi^{*}\in\textbf{FO}^{2}+\textbf{CC}{}$ is obtained from $\Phi$ in the similar manner as $\varphi^{*}$ is constructed from $\varphi$, the new formula will have $p^{*}$ valid cells and, possibly, $p\leq p^{*}$. If we wish to express a complexity bound for $\textbf{C}^{2}$, we should inspect the possible increase in $p$ to obtain $p^{*}$
To deal with an arbitrary $\textbf{C}^{2}$ formula, we need to be able to deal with subformulas such as the one in Sentence 3. As one can observe from both the example above and Algorithm 1, encoding such a sentence in universally quantified $\textbf{FO}^{2}+\textbf{CC}$ requires, in order of appearance, Lemmas 2, 5, 4 and 1. Applying Lemmas 2 and 5 introduces only a constant number of fresh predicates. Hence, the increase in $p$ can be expressed by multiplication with a constant $\beta$. See the proof of Theorem 2 for derivation of a value for $\beta$.
Although the constant $\beta$ may increase our polynomial degree considerably, there is another, much more substantial, influence. An application of Lemma 4, which additionally requires Lemma 1 to deal with unskolemized formulas such as in Formula 22 will introduce $2k$ new predicates. Although $k$ is a parameter of the counting quantifiers, i.e., part of the language, and the language is assumed to be fixed, the size of the encoding of $\textbf{C}^{2}$ in $\textbf{FO}^{2}+\textbf{CC}$ obviously depends on $k$. Hence, the number of cells may also increase with respect to $k$, and, as we formally state below, it in fact does.
@@
For any $m\in\mathds{N}$, there exists a sentence $\Gamma=\varphi\wedge\bigwedge_{i=1}^{m}\left(\forall x\exists^{=k_{i}}y\;\psi_{i}(x,y)\right)$ such that the universally quantified $\textbf{FO}^{2}+\textbf{CC}$ encoding of $\Gamma$ obtained using Lemma 4 has $\mathcal{O}(\prod_{i=1}^{m}\gamma(k_{i})\cdot p)$ valid cells, where $p$ is the number of valid cells of $\varphi$ and $\gamma(k)=(k+2)\cdot 2^{k-1}$.
@@
$\bigwedge_{i=1}^{m}\left(\forall x\exists^{=k_{i}}E_{i}(x,y)\right).$ (24)
In this setting, $\varphi$ is Formula 23 and it has $p=2^{m}$ valid cells. We need to apply Lemma 4 $m$ times to encode sentences in Formula 24 into $\textbf{FO}^{2}+\textbf{CC}$. Let us investigate one such application.
@@
Second, consider valid cells of $\varphi$, that contain $E_{i}(x,x)$ positively. Then exactly one of $f_{ij}(x,x)$ can be satisfied in each cell. That will cause the number of cells to be multiplied by $k_{i}$. Next, for a particular cell, denote $t$ the index such that $f_{it}(x,x)$ is positive in that cell. Then all $A_{ij}(x)$ such that $j\neq t$ will again be free to assume either positive or a negative form. Only $A_{it}(x)$ will be fixed to being positive. Hence, the number of such cells will further be multiplied by $2^{k_{i}-1}$.
Overall, for a single application of Lemma 4, the number of cells will be $\mathcal{O}(2^{k_{i}}\cdot p+k_{i}\cdot 2^{k_{i}-1}\cdot p)\in\mathcal{O}((k_{i}+2)\cdot 2^{k_{i}-1}\cdot p)$, where we upper bound both partitions of the valid cells of $\varphi$ by their total number. After repeated application of Lemmas 4 and 1, the bound above directly leads to the bound we sought to prove. [^3] In the proof, we opted for as simple formula as possible. It would be easy to handle computing WFOMC for $\Gamma$ by decomposing the problem into $m$ identical and independent problems. It is, however, also easy to envision a case where such decomposition is not as trivial. Consider adding constraints such that for each $x$ and $y$, there is only one $i$ such that $E_{i}(x,y)$ is satisfied.
One last consideration for dealing with all of $\textbf{C}^{2}$ is Lemma 3. The lemma adds one new cardinality constraint to the formula, which does not increase number of cells in any way. It will, however, require more calls to an oracle for WFOMC over $\textbf{FO}^{2}$. That influence on the overall bound can still be concentrated in the function $\alpha$. To distinguish the values from before encoding $\textbf{C}^{2}$ and after, let us denote the new value $\alpha^{\prime}$.
@@
where $\Gamma\in\textbf{FO}^{2}+\textbf{CC}$. For any $n\in\mathds{N}$ and any fixed weights $(w,\overline{w})$, $\textup{WFOMC}(\varphi,n,w,\overline{w})$ can be computed in time $\mathcal{O}(n^{\alpha^{\prime}}\cdot n^{1+\prod_{i=1}^{m}\beta\cdot\gamma(k_{i})\cdot p})$, where $p$ is the number of valid cells of $\Gamma$, and $\alpha^{\prime}$ and $\beta$ are constants with respect to both $n$ and each $k_{i}$.
@@
into $\textbf{UFO}^{2}+\textbf{CC}$ using Algorithm 1.
First, we apply Lemma 2 and leverage distributive law obtaining
@@
$\wedge\left(\forall x\;A(x)\Leftrightarrow\left(C(x)\vee D(x)\vee\overline{P_{i}}\right)\right)\wedge\left(\forall x\;A(x)\vee\left(\exists^{=k_{i}}y\;R_{i}(x,y)\right)\right),$
@@
$\mu^{(2)}_{i}$ $=\left(\forall x\;C(x)\vee D(x)\right)\wedge\left(\forall x\;\overline{P_{i}}(x)\Leftrightarrow\neg P_{i}(x)\right)\wedge\left(\forall x\;A(x)\Leftrightarrow\left(C(x)\vee D(x)\vee\overline{P_{i}}\right)\right)$
$\wedge\left(\forall x\forall y\;\neg A(x)\Rightarrow\left(R_{i}(x,y)\Leftrightarrow B^{R_{i}}(x,y)\right)\right)\wedge\left(\forall x\forall y\;\left(A(x)\wedge B^{R_{i}}(x,y)\right)\Rightarrow U^{R_{i}}(y)\right)$
@@
Let us now inspect, how the number of cells may have increased so far. If the original formula had $p$ valid cells, then after application of Lemma 2, there may be up to $4p$ cells. To show that, consider partitioning valid cells into those that satisfy the subformula $\psi(\textbf{x})$ considered by Lemma 2 and those that do not. In the former case, the new cells must satisfy both $C(x)$ and $D(x)$, thus the number of cells remains unchanged. In the latter case, $C(x)$ or $D(x)$ must be satisfied, leading to a factor of three. Hence, we have an upper bound of $p+3p=4p$.
Similarly for Lemma 5, in cells that do not satisfy $A(x)$, $B^{R}(x,x)$ is determined, but $U^{R}(x)$ is unconstrained, leading to a factor of two. In the other cells, if $B^{R}(x,x)$ is also satisfied, then $U^{R}(x)$ is determined, but it is unconstrained if $B^{R}(x,x)$ is satisfied, which can be upper bounded by considering only the second case, leading to another factor of two. Hence, we have another factor of $2+2=4$. Overall, in the worst case, the increase in the number of cells by Lemmas 2 and 5 is by a factor of
@@
$\mu_{i}^{(3)}$ $=\left(\forall x\;C(x)\vee D(x)\right)\wedge\left(\forall x\;\overline{P_{i}}(x)\Leftrightarrow\neg P_{i}(x)\right)\wedge\left(\forall x\;A(x)\Leftrightarrow\left(C(x)\vee D(x)\vee\overline{P_{i}}\right)\right)$
$\wedge\left(\forall x\forall y\;\neg A(x)\Rightarrow\left(R_{i}(x,y)\Leftrightarrow B^{R_{i}}(x,y)\right)\right)\wedge\left(\forall x\forall y\;\left(A(x)\wedge B^{R_{i}}(x,y)\right)\Rightarrow U^{R_{i}}(y)\right)$
@@
$\wedge\left(\forall x\forall y\;B^{R_{i}}(x,y)\Leftrightarrow\bigvee_{i=1}^{k_{i}}f_{i}(x,y)\right)\wedge\bigwedge_{1\leq i<j\leq k_{i}}\left(\forall x\forall y\;f_{i}(x,y)\Rightarrow\neg f_{j}(x,y)\right)$
@@
where $f_{i}/2$ and $A_{i}/1$ are fresh predicates with $\overline{w}(A_{i})=-1$. It follows from Lemma 6 that application of Lemmas 4 and 1 may increase the number of valid cells by a factor of
@@
The factor is an upper bound due to the fact that the sentence $\Gamma$ shown in the proof of Lemma 6 affords the atoms with he predicates $f_{i}$ and $A_{i}$ the highest possible number of degrees of freedom. The truth values of the atoms are determined only by sentences added through application of Lemmas  4 and 1. Therefore, there cannot be more valid cells.
Finally, assuming that each $\mu_{i}$ is encoded into $\textbf{FO}^{2}+\textbf{CC}$ independently of others, we can substitute all values into the bound from Theorem 1 obtaining the bound from our claim.
@@
$\alpha^{\prime}$ $=\left(\sum_{i=1}^{m}arity(U^{R_{i}})+arity(B^{R_{i}})\right)+\left(\sum_{i=1}^{m_{1}}arity(Q_{i})+1\right)=3m+m_{1}+\sum_{i=1}^{m_{1}}arity(Q_{i}).$
@@ 4 Improving the Upper Bound
Let us now inspect the bound from Theorem 2. Although it is polynomial in $n$, meaning that $\textbf{C}^{2}{}$ is in fact domain-liftable (Kuželka, 2021), we can see that the number of valid cells (a part of the polynomial’s degree) grows exponentially with respect to the counting parameters $k_{i}$. In this section, we propose an improved encoding to the one from Lemma 4 which reduces said growth to a quadratic one.
The new encoding does not build on entirely new principles, instead, it takes the existing transformation and makes it more efficient. Same as in Lemma 4, we will describe the situation for dealing with one $\forall\exists^{=k}$ subformula. The procedure could easily be generalized to having $m\in\mathds{N}$ such subformulas by repeating the process for each of them independently.
The largest issue with the current encoding are the Skolemization predicates $A_{i}/1$ which increase the number of valid cells exponentially with respect to $k$. The new encoding will seek to constrain those predicates so that the increase is reduced. Let us start with a formula
@@
where $\varphi\in\textbf{FO}^{2}+\textbf{CC}{}$ and $\psi$ is quantifier-free. Let us also consider the encoding of $\Gamma$ in a universally quantified $\textbf{FO}^{2}+\textbf{CC}$, i.e.,
$\Gamma^{*}$ $=\varphi\wedge\left(\forall x\forall y\;R(x,y)\Leftrightarrow\psi(x,y)\right)\wedge(|R|=k)$

### v2

@@ 3.1 A Worked Example on Removing Counting
with $\overline{w}(D)=-1$. Apart from applying Lemma 2, we also introduced another fresh predicate $\overline{P}/1$, which wraps the negation of $P$ for brevity further down the line.
Next, we need to apply Lemma 5 three times to Sentences 9 through 11. The repetitions can, luckily, be avoided. By the distributive property of conjunctions and disjunctions, we can factor $\overline{P}/1,C/1$ and $D/1$ out from the sentences, thus obtaining
@@
$\wedge\left(\forall x\;F(x)\Leftrightarrow\left(C(x)\wedge D(x)\wedge\overline{P}(x)\right)\right)$ (14)
@@
$\wedge\mathcal{T}_{l}(F,R)\wedge\left(\forall x\;F(x)\Leftrightarrow\left(C(x)\wedge D(x)\wedge\overline{P}(x)\right)\right)$
@@
where $B^{L}/2$ is a fresh predicate. We need Lemma 4 to eliminate this counting construct. Using said lemma will turn $\mathcal{T}_{l}(K,L)$ into
@@
$\wedge\mathcal{T}^{\prime\prime}_{l}(F,R)\wedge\left(\forall x\;F(x)\Leftrightarrow\left(C(x)\wedge D(x)\wedge\overline{P}(x)\right)\right)$
@@ 3.2 Deriving the Upper Bound
As one can observe, the formula $\varphi^{*}$ in the example above has grown considerably compared to its original form. The question is whether the formula growth can influence the asymptotic bound from Theorem 1. At first glance, the answer may seem negative. That is due to the fact that the transformation only extends the vocabulary (which is assumed to be fixed), adds cardinality constraints (which are concentrated in the function $\alpha$), and increases the length of the input formula (which is also constant with respect to $n$). However, there is one caveat to be aware of. When extending the vocabulary, we may introduce new valid cells. The vocabulary is fixed once WFOMC computation starts, but if we have formula $\Gamma\wedge\Phi$ such that $\Gamma\in\textbf{FO}^{2}+\textbf{CC}{}$ and $\Phi\in\textbf{C}^{2}{}$, then this formula already has $p$ valid cells. Once we construct $\Gamma\wedge\Phi^{*}$, where $\Phi^{*}\in\textbf{FO}^{2}+\textbf{CC}{}$ is obtained from $\Phi$ using Algorithm 1, the new formula will have $p^{*}$ valid cells and, possibly, $p\leq p^{*}$. If we wish to express a complexity bound for $\textbf{C}^{2}$, we should inspect the possible increase in $p$ to obtain $p^{*}$.
To deal with an arbitrary $\textbf{C}^{2}$ formula, we need to be able to deal with subformulas such as the one in Sentence 3. As one can observe from both the example above and Algorithm 1, encoding such a sentence in $\textbf{UFO}^{2}+\textbf{CC}$ requires, in order of appearance, Lemmas 2, 5, 4 and 1. Applying Lemmas 2 and 5 introduces only a constant number of fresh predicates. Hence, the increase in $p$ can be expressed by multiplying with a constant $\beta$. See the proof of Theorem 2 for the derivation of a value for $\beta$.
Although the constant $\beta$ may increase our polynomial degree considerably, there is another, much more substantial, influence. An application of Lemma 4, which additionally requires Lemma 1 to deal with unskolemized formulas such as in Formula 22 will introduce $2k$ new predicates. Although $k$ is a parameter of the counting quantifiers, i.e., part of the language, and the language is assumed to be fixed, the size of the encoding of $\textbf{C}^{2}$ in $\textbf{UFO}^{2}+\textbf{CC}$ obviously depends on $k$. Hence, the number of cells may also increase with respect to $k$, and, as we formally state below, it, in fact, does.
@@
For any $m\in\mathds{N}$, there exists a sentence $\Gamma=\varphi\wedge\bigwedge_{i=1}^{m}\left(\forall x\exists^{=k_{i}}y\;\psi_{i}(x,y)\right)$ such that the $\textbf{UFO}^{2}+\textbf{CC}$ encoding of $\Gamma$ obtained using Lemma 4 has $\mathcal{O}(p\cdot\prod_{i=1}^{m}\gamma(k_{i}))$ valid cells, where $p$ is the number of valid cells of $\varphi$ and $\gamma(k)=(k+2)\cdot 2^{k-1}$.
@@
$\bigwedge_{i=1}^{m}\left(\forall x\exists^{=k_{i}}y\;E_{i}(x,y)\right).$ (24)
In this setting, $\varphi$ is Formula 23 and it has $p=2^{m}$ valid cells. We need to apply Lemmas 4 and 1 $m$ times to encode sentences in Formula 24 into $\textbf{UFO}^{2}+\textbf{CC}$. Let us investigate one such application.
@@
Second, consider valid cells of $\varphi$, that contain $E_{i}(x,x)$ positively. Then exactly one of $f_{ij}(x,x)$ can be satisfied in each cell. That will cause the number of cells to be multiplied by $k_{i}$. Next, for a particular cell, denote $t$ the index such that $f_{it}(x,x)$ is positive in that cell. Then all $A_{ij}(x)$ such that $j\neq t$ will again be free to assume either a positive or a negative form. Only $A_{it}(x)$ will be fixed to being positive. Hence, the number of such cells will further be multiplied by $2^{k_{i}-1}$.
Overall, for a single application of Lemma 4, the number of cells will be $\mathcal{O}(p\cdot 2^{k_{i}}+p\cdot k_{i}\cdot 2^{k_{i}-1})\in\mathcal{O}(p\cdot(k_{i}+2)\cdot 2^{k_{i}-1})$, where we upper bound both partitions of the valid cells of $\varphi$ by their total number. After repeated application of Lemmas 4 and 1, the bound above directly leads to the bound we sought to prove. [^4] In the proof, we opted for as simple formula as possible. It would be easy to handle computing WFOMC for $\Gamma$ by decomposing the problem into $m$ identical and independent problems. It is, however, also easy to envision a case where such decomposition is not as trivial. Consider adding constraints such that for each $x$ and $y$, there is only one $i$ such that $E_{i}(x,y)$ is satisfied.
One last consideration for dealing with all of $\textbf{C}^{2}$ is Lemma 3. The lemma adds one new cardinality constraint to the formula, which does not increase the number of cells in any way. It will, however, require more calls to an oracle for WFOMC over $\textbf{FO}^{2}$. That influence on the overall bound can still be concentrated in the function $\alpha$. To distinguish the values from before encoding $\textbf{C}^{2}$ and after, let us denote the new value $\alpha^{\prime}$.
@@
where $\Gamma\in\textbf{FO}^{2}+\textbf{CC}$. For any $n\in\mathds{N}$ and any fixed weights $(w,\overline{w})$, $\textup{WFOMC}(\varphi,n,w,\overline{w})$ can be computed in time $\mathcal{O}(n^{\alpha^{\prime}}\cdot n^{1+p\cdot\prod_{i=1}^{m}\beta\cdot\gamma(k_{i})})$, where $p$ is the number of valid cells of $\Gamma$, $\alpha^{\prime}$ and $\beta$ are constants with respect to both $n$ and the counting parameters $k_{i}$ and $\gamma(k)=(k+2)\cdot 2^{k-1}$.
@@
into $\textbf{FO}^{2}+\textbf{CC}$.
First, following Algorithm 1 and the example in Section 3.1, we apply Lemma 2 and leverage distributive law, obtaining
@@
$\wedge\left(\forall x\;A(x)\Leftrightarrow\left(C(x)\wedge D(x)\wedge\overline{P_{i}}\right)\right)$
$\wedge\left(\forall x\;A(x)\vee\left(\exists^{=k_{i}}y\;R_{i}(x,y)\right)\right),$
@@
$\wedge\left(\forall x\;A(x)\Leftrightarrow\left(C(x)\wedge D(x)\wedge\overline{P_{i}}\right)\right)$
$\wedge\left(\forall x\forall y\;\neg A(x)\Rightarrow\left(R_{i}(x,y)\Leftrightarrow B^{R_{i}}(x,y)\right)\right)$
$\wedge\left(\forall x\forall y\;\left(A(x)\wedge B^{R_{i}}(x,y)\right)\Rightarrow U^{R_{i}}(y)\right)$
@@
Let us now inspect how the number of cells may have increased so far. If the original formula had $p$ valid cells, then after application of Lemma 2, there may be up to $4p$ cells. To show that, consider partitioning valid cells into those that satisfy the subformula $\psi(\textbf{x})$ considered by Lemma 2 and those that do not. In the former case, the new cells must satisfy both $C(x)$ and $D(x)$. Thus, the number of cells remains unchanged. In the latter case, $C(x)$ or $D(x)$ must be satisfied, leading to a factor of three. Hence, we have an upper bound of $p+3p=4p$.
Similarly, for Lemma 5, in cells that do not satisfy $A(x)$, $B^{R}(x,x)$ is determined, but $U^{R}(x)$ is unconstrained, leading to a factor of two. In the other cells, if $B^{R}(x,x)$ is also satisfied, then $U^{R}(x)$ is determined. However, it is unconstrained if $B^{R}(x,x)$ is satisfied, which can be upper bounded by considering only the second case, leading to another factor of two. Hence, we have another factor of $2+2=4$. Overall, in the worst case, the increase in the number of cells by Lemmas 2 and 5 is by a factor of
@@
$\mu_{i}^{(3)}$ $=\left(\forall x\;C(x)\vee D(x)\right)\wedge\left(\forall x\;\overline{P_{i}}(x)\Leftrightarrow\neg P_{i}(x)\right)$
$\wedge\left(\forall x\;A(x)\Leftrightarrow\left(C(x)\wedge D(x)\wedge\overline{P_{i}}\right)\right)$
$\wedge\left(\forall x\forall y\;\neg A(x)\Rightarrow\left(R_{i}(x,y)\Leftrightarrow B^{R_{i}}(x,y)\right)\right)$
$\wedge\left(\forall x\forall y\;\left(A(x)\wedge B^{R_{i}}(x,y)\right)\Rightarrow U^{R_{i}}(y)\right)$
@@
$\wedge\left(\forall x\forall y\;B^{L}(x,y)\Leftrightarrow\bigvee_{i=1}^{k_{i}}f_{i}(x,y)\right)$
$\wedge\bigwedge_{1\leq i<j\leq k_{i}}\left(\forall x\forall y\;f_{i}(x,y)\Rightarrow\neg f_{j}(x,y)\right)$
@@
where $f_{i}/2$ and $A_{i}/1$ are fresh predicates with $\overline{w}(A_{i})=-1$. It follows from Lemma 6 that applying Lemmas 4 and 1 may increase the number of valid cells by a factor of
@@
The factor is an upper bound because the sentence $\Gamma$ shown in the proof of Lemma 6 affords the atoms with the predicates $f_{i}$ and $A_{i}$ the highest possible number of degrees of freedom. The truth values of the atoms are determined only by sentences added through the application of Lemmas  4 and 1. Therefore, there cannot be more valid cells.
Finally, assuming that each $\mu_{i}$ is encoded into $\textbf{FO}^{2}+\textbf{CC}$ independently of others, we can substitute all values into the bound from Theorem 1, obtaining the bound from our claim.
@@
$\alpha^{\prime}$ $=\left(\sum_{i=1}^{m}arity(U^{R_{i}})+arity(B^{R_{i}})\right)$
$\hskip 28.45274pt+\left(\sum_{i=1}^{m_{1}}arity(Q_{i})+1\right)$
$=3m+m_{1}+\sum_{i=1}^{m_{1}}arity(Q_{i}).$
@@ 4 Improving the Upper Bound
Let us now inspect the bound from Theorem 2. Although it is polynomial in $n$, meaning that $\textbf{C}^{2}{}$ is, in fact, domain-liftable (Kuželka, 2021), we can see that the number of valid cells (a part of the polynomial’s degree) grows exponentially with respect to the counting parameters $k_{i}$. In this section, we propose an improved encoding to the one from Lemma 4 which reduces said growth to a quadratic one.
The new encoding does not build on entirely new principles, instead, it takes the existing transformation and makes it more efficient. As in Lemma 4, we will describe the situation for dealing with one $\forall\exists^{=k}$-quantified subformula. The procedure could easily be generalized to having $m\in\mathds{N}$ such subformulas by repeating the process for each of them independently.
The most significant issue with the current encoding are the Skolemization predicates $A_{i}/1$, which increase the number of valid cells exponentially with respect to $k$. The new encoding will seek to constrain those predicates so that the increase is reduced. Let us start with a formula
@@
where $\varphi\in\textbf{FO}^{2}+\textbf{CC}{}$ and $\psi$ is quantifier-free. Let us also consider the encoding of $\Gamma$ in $\textbf{UFO}^{2}+\textbf{CC}$, i.e.,
$\Gamma^{*}$ $=\varphi\wedge\left(\forall x\forall y\;R(x,y)\Leftrightarrow\psi(x,y)\right)\wedge(|R|=n\cdot k)$